OpenStudy (anonymous):

If someone could guide me through the process of solving this integration problem that would be very helpful.

OpenStudy (anonymous):

It's an applications problem dealing with the volume generated by rotating a curve by an axis. I'll go strait to the problem: I have to rotate the region bounded by the graphs y=secx, y=1, x=-1, and x=1 with reference to the x-axis.

OpenStudy (anonymous):

The integral I came with is:\[\int\limits_{-1}^{1}((1+\sec(x))^2-1)dx\] Before proceeding is the integrand correct?

OpenStudy (turingtest):

always graph it first...

OpenStudy (anonymous):

Thanks for the advice but I think I already graphed it and I'm struggling with the integral the graph suggests to be the answer for this problem. I don't think we have covered the indefinite integral of sec(x) in our class.

OpenStudy (turingtest):

ok, I just have to do it myself to make sure you got the right setup I can't just "see" the answer lol, still human...

OpenStudy (turingtest):

the way I see it the outer radius of each disk is \(\sec x\) and the inner is \(y=1\) going about x this gives the integral\[\pi\int \sec^2x-1dx\]so I think you had it wrong...

OpenStudy (turingtest):

this integral really should not be a problem; you should know the integral of sec^2xdx

OpenStudy (turingtest):

bounds are -1 to 1 of course...

OpenStudy (anonymous):

Did you take into account the distance between secx and y=0?

OpenStudy (turingtest):

yes, that is the outer radius\[r_o=y=\sec x\]

OpenStudy (turingtest):

the inner radius is\[r_i=y=1\]

OpenStudy (anonymous):

I think I see how you approach this, I saw the outer disk as secx+1 and the inner disk as 1. I mean their radii.

OpenStudy (turingtest):

you set up like you were revolving around the line y=-1

OpenStudy (turingtest):

ok, not quite, but almost like that you sort of mixed the two ideas...

OpenStudy (anonymous):

Yeah I see what you mean, I was thrown off by the graph of secx and y=1 so close together.

OpenStudy (anonymous):

I really appreciate your help and time. Thanks for everything

OpenStudy (turingtest):

which is why I always say to draw the graph do it by hand, then check with wolfram or a graphing calculator to be sure when you can visualization is the key to these problems

OpenStudy (anonymous):

One last question though, wouldn't the inner disk be a cylinder?

OpenStudy (turingtest):

yes, here it is with a few rings drawn in